OpenStudy (anonymous):

how much is .50 cent per minute? And how would that be for an hour?

OpenStudy (anonymous):

there are 60min in an hour right

OpenStudy (unklerhaukus):

\[0.50\left[\frac ¢{\text{min}}\right]\]\[=0.50\left[\frac ¢{\text{min}}\right]\times\left[\frac ${100¢}\right]\]\[=\frac{0.50}{100}\left[\frac ${\text{min}}\right]\]

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

.5 cents is half a penny so first calculate how many cents you have in an hour 1cent*60

OpenStudy (anonymous):

after you get that answer divide by 2 to receive cents/hr

OpenStudy (anonymous):

i find it easier to save fractions till the end of the equation

OpenStudy (anonymous):

so it is 30

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

remember to put your units

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

np do you understand how i did it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

hero (hero):

\[\frac{.50}{\min} = \frac{x}{\text{hr}}\] \[\frac{.50}{\min} = \frac{x}{60\min}\] \[.50 \times 60\min = x \min\] \[\frac{30 \min}{\min} = \frac{x\min}{min}\] \[30 = x\]